If 𝒍=πŸπŸŽπŸŽπ’Žπ’Ž and π’˜=πŸπŸ“πŸŽπ’Žπ’Ž calculate:

a) The value of x which will give the maximum volume

b) The maximum volume of the box

c) Comment of the value obtained in part b.


1
Expert's answer
2021-10-26T08:58:53-0400

Since x is the side of the square that has to be cutted from each corner in order to make a box, then the volume of the box can be expressed in terms of x following way:

"V(x) = (l-2x)*(w-2x)*x = (200-2x)*(150-2x)*x"

The point is to find such value of x that maximizing V(x). We should find derivative of the V(x)

"V'(x) = 3000-800x-600x+12x^{2}=12x^{2}-1400x+30000" and find such x that V'(x) = 0

"12x^{2}-1400x+30000=0\to 3x^{2} - 350x+7500=0"

After solving this equation we received "x{\scriptscriptstyle 1}=28.29mm, x{\scriptscriptstyle 2}=88.38mm". Second root doesn't satisfy the conditions of the task, cause x cannot be greater than "{\frac w 2} = 75". In point "x{\scriptscriptstyle 1}" V'(x) change sign from + to -, so it's a point of maximum of V(x).

"V(x{\scriptscriptstyle 1})=379037.81mm^{2}"

So, we find out that if you have a rectangular with sizes 150mmX200mm(or 0.15mX0.2 m) you can make a box with the max roomines of 379037.81"mm^{2}" (or approximately 0.379"m^{2}")